Use the following information for the Exercises below.

[The following information applies to the questions displayed below.]

Hart Company made 4,900 bookshelves using 22,000 board feet of wood costing $261,800. The company's direct materials standards for one bookshelf are 5 board feet of wood at $11.80 per board foot.

Exercise 23-13 Computation and interpretation of materials variances LO P2

AQ = Actual Quantity
SQ = Standard Quantity
AP = Actual Price
SP = Standard Price

(1) Compute the direct materials price and quantity variances incurred in manufacturing these bookshelves.

Actual Cost 0 0 Standard Cost
0
$0 0 $0
$0
$0
Answer #1
Material price variance = (SP-AP)*AQ
material price variance = (11.80-11.90)*22000
material price variance = 2200 Unfavorable
material quantity variance = (SQ-AQ)*SP
Material quantity variance = (5*4900-22000)*11.80
material quantity variance = 29500 Favorable